OpenStudy (anonymous):

Solve the formula for the indicated variable. P=b+a+3c, for b b=

OpenStudy (ghazi):

P=b+a+3c subtract a and 3c from both the sides P-a-3c=b

OpenStudy (anonymous):

oh ok so B= P-a-3c

OpenStudy (ghazi):

yes b=P-a-3c

OpenStudy (anonymous):

can you help me with another problem please.

OpenStudy (ghazi):

i would love to but i have to go for now, but later, for sure :)

OpenStudy (anonymous):

ok Ghazi

OpenStudy (ghazi):

:)

OpenStudy (ghazi):

there are lot of people they can help you don't worry keep shooting :)

OpenStudy (anonymous):

k